### v2

# Universal Acceleration of Quantum Relaxation by Reset-Induced Mpemba Effect

Abstract
Slow relaxation processes spanning widely separated timescales pose fundamental challenges for probing steady-state properties and engineering functional quantum systems, such as quantum heat engines and quantum computing devices. We introduce a universal quantum reset protocol that enables exponential acceleration of relaxation in general Markovian open quantum systems. The protocol induces a quantum Mpemba effect, wherein systems initialized farther from equilibrium can relax faster than those closer to it, and crucially, this acceleration persists even when the slowest decaying Lindbladian modes form complex-conjugate pairs. Unlike previous approaches, which typically target a single mode, our protocol eliminates multiple relaxation modes simultaneously. This framework provides a versatile and experimentally feasible tool for controlling quantum relaxation timescales, with broad implications for quantum thermodynamics, computation, and state preparation.
Introduction.—Relaxation processes in open quantum systems, wherein a driven or interacting system settles into an equilibrium or a nonequilibrium steady state, are of fundamental importance in nonequilibrium physics and for practical quantum technologies. The timescales of such relaxations often determine the feasibility and performance of quantum devices. For example, in quantum thermodynamics, the power output of heat engines or transport junctions is limited by how rapidly the working substance reaches its steady state in each cycle [1, 2, 3]. In addition, fast relaxation facilitates ground state laser cooling [4, 5, 6] and the reliable preparation of quantum states [7]. Faster relaxation can also lead to more efficient quantum algorithms [8], particularly in dissipative computational architectures where the output is encoded in the system’s stationary state [8, 9]. Slow relaxation inevitably allows other unwanted dissipative dynamics to consume substantial time resources, compromising the efficiency of the final stabilized state [10, 11]. In contrast, slow relaxation is sometimes desirable—such as when the computational task targets metastable states [12, 13, 14, 15], whose prolonged lifetimes are essential for robust information storage or approximate optimization.
Researchers have developed various techniques to accelerate relaxation toward both thermal equilibrium [16] and nonequilibrium steady states [17, 18, 11]. Most of these methods, however, are restricted to specific systems and rely on intricate external controls. For instance, some approaches require auxiliary Hamiltonians or full knowledge of the system Hamiltonian, limiting their applicability to simple setups. A more general and elegant strategy is to accelerate relaxation by tailoring the initial state distribution, without the need for continuous external driving. A particularly intriguing example of this approach is inspired by the so-called Mpemba effect [19], a counterintuitive relaxation anomaly in which a hotter system can cool faster than a colder one. This phenomenon implies the existence of an optimal initial state that minimizes the relaxation timescale. Theoretical frameworks for identifying such optimal states have been proposed in both classical [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and quantum systems [10, 32, 33, 34, 35, 36, 37, 38, 39], see also [40, 41] for recent reviews. However, existing quantum implementations typically impose unrealistic constraints, such as assuming pure initial states or requiring the second-largest eigenvalue of the Lindbladian to be real, which limit their applicability in general settings. Thus, a general and practical method for inducing faster relaxation toward arbitrary steady states in open quantum systems remains elusive. Likewise, a systematic recipe for prolonging the lifetime of metastable states, i.e., for slowing down the relaxation process is also lacking.
In this study, we address these challenges by proposing an experimentally feasible protocol that can both accelerate and decelerate general quantum relaxation processes. The protocol is based on the quantum reset operation [42, 43, 44, 45, 46], which is a key component of many quantum computation tasks [47]. For instance, qubit initialization to the computational ground state $|00...0\rangle$ is typically realized through a reset procedure. Recent experimental advancements have demonstrated the practical implementation of quantum reset, e.g. Aamir et al. have realized an autonomous quantum absorption refrigerator using superconducting circuits, capable of resetting a transmon qubit [48]. Other reset schemes have also been proposed [49, 50, 51, 52]. In our protocol, we introduce a finite-duration reset phase into the open-system dynamics by temporarily coupling the system to a reset channel. We show that this operation can exponentially accelerate relaxation, even when starting from mixed initial states and in the presence of complex decay modes. Remarkably, our method allows for the simultaneous elimination of multiple relaxation modes—a capability absent in previous approaches. Moreover, by applying the protocol selectively, the lifetime of metastable states can be extended, offering a systematic route to suppress relaxation.
Setup.—We consider a general open quantum system defined on a Hilbert space $\mathcal{H}$ of dimension $d$, whose dynamics in the interaction picture are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) master equation $d\rho/dt=\mathcal{L}(\rho),$ where the Lindbladian superoperator $\mathcal{L}$ is given by
$\mathcal{L}(\rho)\equiv-i[H,\rho]+\sum_{i}\left[J_{i}\rho J_{i}^{\dagger}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},\rho\}\right].$ (1)
[…]
$\text{Tr}(L_{k}^{\dagger}R_{h})=\delta_{kh}.$ (2)
The dual superoperator $\mathcal{L}^{\dagger}$ governs the Heisenberg-picture dynamics of observables:
$\mathcal{L}^{\dagger}(O)=i[H,O]+\sum_{i}\left[J_{i}OJ_{i}^{\dagger}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},O\}\right].$
Given an initial state $\rho_{0}$, the system state at time $t$ evolves as
$\rho(t)=e^{t\mathcal{L}}[\rho_{0}]=\rho_{ss}+\sum_{k=2}^{d^{2}}e^{\lambda_{k}t}c_{k}R_{k},$ (3)
where $c_{k}\equiv\text{Tr}(L_{k}^{\dagger}\rho_{0})$. The unique stationary state of the open quantum system $\rho_{ss}$ is given by $\rho_{ss}=\lim_{t\rightarrow\infty}\rho(t)=R_{1}$, assuming $\lambda_{1}=0$ is non-degenerate. In the long-time limit, the relaxation is dominated by the slowest decaying mode, and the deviation from stationarity obeys
$||\rho(t)-\rho_{ss}||\sim\exp\left(|\text{Re}\lambda_{2}|t\right),$
so the characteristic relaxation timescale is given by $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$. For clarity of exposition, we first consider the simple case where the second eigenvalue $\lambda_{2}$ is real and non-degenerate, as in [10]. Later, we will generalize the discussion to include cases where the subleading eigenvalues form complex conjugate pairs.
Reset protocol to accelerate relaxation processes.—Recognizing that the slowest decaying mode governs the relaxation timescale, we propose a quantum reset protocol that suppresses its excitation from a general initial state via a finite-duration reset phase, thereby enabling exponential acceleration of relaxation. Specifically, we let the system evolve under modified dynamics for a finite time $t_{s}$, during which it is stochastically reset to a chosen state $|\delta\rangle$ at random times, with events following a Poisson process of rate $r$. The reset state may also be taken as a mixed state $\rho_{\delta}$. After time $t_{s}$, the reset channel is turned off, and the system resumes evolving under the original Lindbladian $\mathcal{L}$. The modified dynamics during the reset phase is governed by a new Lindbladian $\mathcal{L}_{r}$, defined as
$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)|\delta\rangle\langle\delta|-r\rho.$ (4)
This can be physically implemented by adding reset jump operators [42] $J_{i}^{r}=\sqrt{r}|\delta\rangle\langle\phi_{i}|$ to the Lindbladian (1), where $|\phi_{i}\rangle$ form an auxiliary complete orthonormal basis. When the reset state is a mixed state $\rho_{\delta}$, the new Lindbladian reads $\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)\rho_{\delta}-r\rho$. Importantly, this protocol modifies the relaxation spectrum without altering the eigenmodes. This is because the term $r\text{Tr}(\rho)|\delta\rangle\langle\delta|$ in the modified Lindbladian contributes only to the stationary mode and vanishes on all traceless eigenmodes $R_{i}$ with $i\geq 2$, which satisfy $\mathrm{Tr}(R_{i})=0$ due to $\mathrm{Tr}(L_{1}^{\dagger}R_{i})=0$ and $L_{1}^{\dagger}=\mathbb{I}$ [cf. (2)]. As a result, the modified Lindbladian acts as
$\mathcal{L}_{r}(R_{i})=(\lambda_{i}-r)R_{i},\ i\in\{2,...,d^{2}\},$ (5)
leading to a uniform spectral shift by $-r$ while preserving the mode structure. Let $\rho^{r}(t):=e^{t\mathcal{L}_{r}}\rho_{0}$ be the state at time $t$ under reset dynamics. Its spectral decomposition reads:
$\rho^{r}(t)=$ $e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})e^{(\lambda_{k}-r)t}R_{k}$
::=:= $\rho_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}R_{k},$ (6)
where $c_{k}^{r}(t)\propto\text{Tr[$L_{k}^{\dagger}$$\rho^{r}(t)$]}$ are time-dependent mode amplitudes. Explicitly, these modified coefficients are given by
$c_{k}^{r}(t)\equiv\left[c_{k}-\frac{r\cdot d_{k}}{r-\lambda_{k}}\right]e^{-rt}+\frac{r\cdot d_{k}}{r-\lambda_{k}}e^{-\lambda_{k}t},$ (7)
where $d_{k}=\langle\delta|L_{k}^{\dagger}|\delta\rangle$. By properly choosing $t_{s}$, we can set $c_{2}^{r}(t_{s})=0$, thereby eliminating the overlap with the slowest decaying mode. Solving $c_{2}^{r}(t_{s})=0$ yields
$t_{s}(r)=\frac{1}{r-\lambda_{2}}\ln\left[1-\frac{c_{2}(r-\lambda_{2})}{r\cdot d_{2}}\right].$ (8)
This condition is always achievable by an appropriate choice of reset state $|\delta\rangle$ ensuring $t_{s}(r)\geq 0$. After the reset phase ($t>t_{s}$), the system returns to original Lindbladian evolution, now without the contribution from the slowest mode. The full evolution reads:
$\rho^{r}(t)=$ $\rho_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t)R_{k}$
$+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}\Theta(t-t_{s})R_{k},$ (9)
where $\Theta(x)$ is the Heavside step function. This protocol reduces the relaxation time to
$\tau_{{\rm rel}}\sim t_{s}+\frac{1}{|{\rm Re}(\lambda_{3})|},$ (10)
which is generically shorter than the unperturbed timescale $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$.
Our protocol also enables controlled deceleration of relaxation. By choosing $|\delta\rangle$ and $t_{s}$ to amplify the magnitude of $c_{k}^{r}(t)$, one can extend the lifetime of selected metastable modes. This feature is especially useful in quantum metrology and computation, where long-lived metastable states play a functional role. Our reset protocol thus provides a general framework for controlling relaxation processes in open quantum systems.
Extension to complex-conjugate relaxation modes.—We now consider the more general case where the dominant relaxation modes form a complex-conjugate pair. In this case, the long-time dynamics of the density matrix takes the form
$\rho(t)=\rho_{ss}+e^{\lambda_{2}t}c_{2}R_{2}+e^{\lambda_{2}^{\dagger}t}c_{2}^{*}R_{2}^{\dagger}+...,$ (11)
where $c_{2}=\text{Tr}(L_{2}^{\dagger}\rho_{0})$ and $c_{2}^{*}=\text{Tr}(L_{2}\rho_{0})$ is its complex conjugate. Most known strategies are unable to simultaneously suppress both components of such complex relaxation modes.
Our reset protocol, however, naturally addresses this limitation. The modified coefficients under quantum reset are given by
$c_{2}^{r}(t)=\left[c_{2}-\frac{r\cdot d_{2}}{r-\lambda_{2}}\right]e^{-rt}+\frac{r\cdot d_{2}}{r-\lambda_{2}}e^{-\lambda_{2}t}$ ,
$c_{2}^{r,*}(t)=\left[c_{2}^{*}-\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}\right]e^{-rt}+\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}e^{-\lambda_{2}^{*}t}$
Importantly, we observe that these coefficients remain complex conjugates of each other for all $t$, i.e., $[c_{2}^{r}(t)]^{*}=c_{2}^{r,*}(t)$. Therefore, eliminating the entire complex mode reduces to a single complex condition:
$c_{2}^{r}(t_{s})=0\Leftrightarrow\text{Re}[c_{2}^{r}(t_{s})]=\text{Im}[c_{2}^{r}(t_{s})]=0,$ (12)
because $c_{2}^{r}(t_{s})=0\Rightarrow c_{2}^{r,*}(t_{s})=[c_{2}^{r}(t_{s})]^{*}=0$. This condition can be fulfilled by tuning the reset rate $r$ and the reset duration $t_{s}$, thereby simultaneously eliminating both components of the dominant complex relaxation mode. In practice, it suffices to reduce $|c_{2}^{r}(t_{s})|$ to an exponentially small value ($\text{Re}[c_{2}^{r}(t_{s})]\ll 1,\ \text{Im}[c_{2}^{r}(t_{s})]\ll 1$), which still leads to exponential acceleration of relaxation. We demonstrate this behavior concretely in the following example.
Example.—To illustrate our protocol, we consider a minimal example: a two-level quantum system, which may represent a qubit or a spin-$1/2$ particle. We set the ground state energy $E_{0}=0$ and excited state energy $E_{1}=E$, such that the system Hamiltonian reads
$H=E|1\rangle\langle 1|+\Omega(|0\rangle\langle 1|+|1\rangle\langle 0|)\text{,}$
where $\Omega$ is the intrinsic coherent coupling between the two levels. The system is coupled to a thermal environment through the Lindblad jump operators $J_{0}=|1\rangle\langle 0|=\sigma_{+}$ and $J_{1}=|0\rangle\langle 1|=\sigma_{-}$. The resulting open-system dynamics are governed by the master equation:
$\frac{d\rho(t)}{dt}=$ $-i[H,\rho]+\Gamma_{0}\left(J_{0}\rho J_{1}-\frac{1}{2}\{J_{0}J_{1},\rho\}\right)$
$+\Gamma_{1}\left(J_{1}\rho J_{0}-\frac{1}{2}\{J_{1}J_{0},\rho\}\right),$ (13)
where $\Gamma_{0,1}$ are the transition rates from state $|0\rangle$ to state $|1\rangle$ or converse. Assuming the rates satisfy detailed balance with inverse temperature $\Gamma_{0}e^{-\beta_{\text{env}}E_{0}}=\Gamma_{1}e^{-\beta_{\text{env}}E_{1}}$, i.e., $\Gamma_{0}=\Gamma_{1}e^{-\beta_{\text{env}}E},$ the system evolves to a unique equilibrium state $\rho_{eq}=\frac{e^{-\beta_{\text{env}}H}}{\text{Tr}(e^{-\beta_{\text{env}}H})}$ irrespective of the initial condition. We initialize the system in a general mixed state:
$\hat{\rho}_{0}$ $=\left(\begin{array}[]{cc}\frac{1}{1+e^{-\beta_{0}E}}&ke^{i\phi}\\ ke^{-i\phi}&\frac{e^{-\beta_{0}E}}{1+e^{-\beta_{0}E}}\end{array}\right),$
$0\leq k$ $\leq\frac{\sqrt{e^{\beta_{0}E}}}{e^{\beta_{0}E}+1},\quad 0\leq\phi\leq 2\pi,$
where the off-diagonal terms $ke^{\pm i\phi}$ quantify initial quantum coherence. In the absence of intrinsic transitions ($\Omega=0$), the Hamiltonian is diagonal and the relaxation process reduces to pure decoherence. In this case, all eigenmodes and eigenvalues of the Lindbladian can be computed analytically (see Appendix B). When we apply our reset protocol with the reset state $|\delta\rangle=|0\rangle\langle 0|$, corresponding to the ground state to ensure $t_{s}(r)\geq 0$, the two slowest decaying modes form a complex-conjugate pair:
$c_{2}^{r}(t)=e^{-rt+i\phi}k,\quad c_{2}^{r,*}(t)=e^{-rt-i\phi}k.$ (14)
This overlap cannot be entirely eliminated in finite time, but it decays exponentially fast, leading to a relaxation timescale
$\tau_{{\rm rel}}\sim\frac{1}{r}+\frac{1}{|\text{Re}(\lambda_{3})|},$ (15)
where $1/r$ characterizes the reset protocol timescale, and $1/|\text{Re}(\lambda_{3})|$ represents the accelerated relaxation timescale with the slowest mode eliminated. Clearly, $\tau_{{\rm rel}}<1/|{\rm Re}(\lambda_{2})|$ for sufficiently large $r$. This structure is general and holds for more complex systems due to analogous spectral modifications introduced by the quantum reset. By appropriately tuning the reset rate, the relaxation timescale can approach $1/|\rm Re(\lambda_{3})|$ arbitrarily closely. Using Eq. (9), the full dynamics under our reset protocol can thus be determined.
Since the condition $c_{2}^{r}(t_{s})=0$ cannot be satisfied exactly in this case—though $|c_{2}^{r}(t_{s})|$ can be made exponentially small by choosing $t_{s}\sim 1/r$—we instead define $t_{s}$ such that a faster decaying mode is suppressed. This pragmatic approach effectively facilitates experimental implementation. Specifically, solving $c_{4}^{r}(t_{s})=0$ gives
$t_{s}=\frac{\ln\left[1+\frac{\Gamma_{1}k_{\text{tot}}}{r}\left(\frac{1}{\Gamma_{0}+\Gamma_{1}e^{(\beta_{0}-\beta_{\text{env}})E}}-\frac{1}{\Gamma_{0}+\Gamma_{1}}\right)\right]}{k_{\text{tot}}},$ (16)
where the total transition rate is $k_{\text{tot}}=r+\Gamma_{0}+\Gamma_{1}$. This choice eliminates the fastest decaying relaxation mode, ensuring the slowest mode becomes exponentially negligible for $t_{s}>1/r$. Importantly, alternative selections for $t_{s}$ can also achieve accelerated relaxation, indicating that detailed dynamics knowledge is not essential. Notably, if the initial inverse temperature satisfies $\beta_{0}<\beta_{\text{env}}$, a scenario analogous to “cooling”, the ground state $|0\rangle\langle 0|$ universally serves as an effective reset state regardless of the reset rate $r$. Conversely, when $\beta_{0}>\beta_{\text{env}}$, resetting to the ground state slows relaxation, necessitating instead the excited state $|1\rangle\langle 1|$ as the reset state for acceleration. This example demonstrates that reset-induced acceleration or deceleration of relaxation is readily controllable through appropriate state selection.
To illustrate the effect of the reset protocol, we next present numerical results. We characterize the distance between the transient state $\rho(t)$ and the stationary state $\rho_{ss}$ using the standard quantum trace distance
[…]
where $|A|:=\sqrt{A^{\dagger}A}$. Other measures, such as the L-infinity norm, yield qualitatively similar outcomes. Utilizing this metric, we illustrate the relaxation dynamics by plotting trace distances from $\rho(t)$ to equilibrium $\rho_{\text{eq}}$ over time $t$ for various rate $r$ in Figure 1 (a). The reset scenario employs initial conditions $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$, whereas the non-reset scenario has $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$. With the environment at lower temperature $\beta_{\text{env}}=4.0$, the non-reset system starts closer to equilibrium. Nonetheless, strikingly, the reset protocol substantially accelerates relaxation, manifesting an anomalous effect reminiscent of the quantum Mpemba effect—systems initially farther from stationary state reach stationarity faster. We next plot the critical time $t_{s}$ required to close the reset as a function of the resetting rate $r$ in Figure 1 (b). The results demonstrate that $t_{s}$ decreases monotonically as $r$ increases, with all other parameters held constant. This finding indicates that reset protocols employing larger resetting rates facilitate more rapid relaxation of open quantum systems.
Figure 1: Relaxation dynamics of the two-state system. (a) and (c): Distances between $\rho(t)$ and $\rho_{\text{eq}}$ as the function of time $t$ with different resetting rate $r$. (b) and (d): The $t_{s}$ as a function of the resetting rate $r$. The parameters are chosen as $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$ with reset protocol and $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$ without reset. For (a) and (b) $\Omega=0$, for (c) and (d) $\Omega=2.0$.
When intrinsic transitions exist ($\Omega\neq 0$), the reset protocol remains effective. In Figure 1 (c) confirms accelerated relaxation for $\Omega=2$, here measured by the L-infinity norm to reduce fluctuations in the curves.
Accelerating relaxation of quantum observables.—In experimental settings, observables rather than full density matrices are typically measured. Our protocol naturally extends to controlling the relaxation dynamics of arbitrary quantum observables. The time evolution of an observable $O$ under unperturbed dynamics reads:
$\langle O\rangle(t)={\rm Tr}[Oe^{\mathcal{L}t}(\rho_{0})]=\langle O\rangle_{ss}+\sum_{k=2}^{d^{2}}c_{k}e^{\lambda_{k}t}{\rm Tr}[OR_{k}].$ (17)
Applying our reset protocol, the modified observable dynamics becomes
$\langle O\rangle^{r}(t)$ $=\langle O\rangle_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t){\rm Tr}[OR_{k}]$
$+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}\Theta(t-t_{s}){\rm Tr}[OR_{k}].$ (18)
Thus, the same reset condition $c_{2}^{r}(t_{s})\approx 0$ also ensures fast convergence of $\langle O\rangle(t)$ to its steady-state value $\langle O\rangle_{ss}$. This allows for direct experimental application of our protocol at the level of observables, without requiring full state tomography.
Concluding remarks.—We have introduced a general framework for accelerating relaxation in open quantum systems via quantum reset dynamics, applicable to arbitrary initial states. The proposed protocol is simple to implement: it requires temporarily coupling the system to a reset channel and turning it off after a finite duration $t_{s}$, which can be tuned through the reset rate $r$. Additionally, the protocol can suppress multiple relaxation modes simultaneously, enabling enhanced control over the relaxation dynamics. In the future study, our approach could be extended to certain non-Markovian and periodically driven (Floquet) dissipative systems, and may offer a route to circumvent known trade-offs in relaxation control [18]. Overall, our results establish quantum reset as a versatile and experimentally feasible tool for controlling relaxation timescales in open quantum dynamics.
This work is partly supported by the Innovation Program for Quantum Science and Technology (2021ZD0303306), MOST(2022YFA1303100).
Appendix
Appendix A Derivation of the modified lindbladian
Acting the time-evolution operator generated by this generalized Lindblad operator on the eigenvalue-expanded form of the initial state
$\rho_{0}=\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})R_{k}$ (19)
results in
$\rho^{r}(t)$ $=e^{t\mathcal{L}_{r}}\rho_{0}$
$=e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})e^{(\lambda_{k}-r)t}R_{k}.$ (20)
To get the detailed form of the first term, note that the stationary state under reset dynamics is given by
$\rho_{ss}^{r}=\lim_{t\rightarrow\infty}\rho^{r}(t)=\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}R_{k},$ (21)
and the stationary state has the property that
$e^{t\mathcal{L}_{r}}\rho_{ss}^{r}=\rho_{ss}^{r}.$ (22)
Thus, acting $e^{t\mathcal{L}_{r}}$ on both sides of the Eq.(21) and combining Eq.(22) bring about
$e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}e^{(\lambda_{k}-r)t}R_{k}$
$=\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}R_{k}$ (23)
Appendix B Spectral analysis of the two-state model
The Eq. (13) can be rewritten as a matrix equation for the vector $\vec{\rho}(t)=[\rho_{00}(t),\rho_{01}(t),\rho_{10}(t),\rho_{11}(t)]^{T}$:
$\frac{d\vec{\rho}(t)}{dt}=L\vec{\rho}(t),$ (24)
where $L$ is a $4\times 4$ Liouvillian matrix in the Fock-Liouvillian space, taking the form as
$L=\left(\begin{array}[]{cccc}-\Gamma_{0}&-i\Omega&i\Omega&\Gamma_{1}\\ -i\Omega&-iE-\frac{\Gamma_{0}+\Gamma_{1}}{2}&0&i\Omega\\ i\Omega&0&iE-\frac{\Gamma_{0}+\Gamma_{1}}{2}&-i\Omega\\ \Gamma_{0}&i\Omega&-i\Omega&-\Gamma_{1}\end{array}\right).$ (25)
In the simple case $\Omega=0$, the right eigenvalues and eigenvectors ot the matrix $L$ are given by
$\lambda_{1}$ $=0,\ \vec{R}_{1}=(\Gamma_{1},0,0,\Gamma_{0})^{\text{T}}$
$\lambda_{4}$ $=-\Gamma_{0}-\Gamma_{1},\ \vec{R}_{2}=(-1,0,0,1)^{\text{T}}$
$\lambda_{2}$ $=\frac{1}{2}(-2i-\Gamma_{0}-\Gamma_{1}),\ \vec{R}_{3}=(0,1,0,0)^{\text{T}}$
$\lambda_{3}$ $=\frac{1}{2}(2i-\Gamma_{0}-\Gamma_{1}),\ \vec{R}_{4}=(0,0,1,0)^{\text{T}}$ (26)
Likewise, the left eigenvalues and eigenvectors are given by
$\lambda_{1}^{\star}$ $=0,\ \vec{L}_{1}=(1,0,0,1)^{\text{T}}$
$\lambda_{4}^{\star}$ $=-\Gamma_{0}-\Gamma_{1},\ \vec{L}_{2}=(-\Gamma_{0},0,0,\Gamma_{1})^{\text{T}}$
$\lambda_{2}^{\star}$ $=\frac{1}{2}(2i-\Gamma_{0}-\Gamma_{1}),\ \vec{L}_{3}=(0,1,0,0)^{\text{T}}$
$\lambda_{3}^{\star}$ $=\frac{1}{2}(-2i-\Gamma_{0}-\Gamma_{1}),\ \vec{L}_{4}=(0,0,1,0)^{\text{T}}.$ (27)
Then, by Eq.(8) in the main text, the expression for the critical time $t_{s}$ is given by
$t_{s}=\frac{\ln\left[1+\frac{\Gamma_{1}(r+\Gamma_{0}+\Gamma_{1})\left(\frac{1}{\Gamma_{0}+\Gamma_{0}e^{\beta_{0}E}}-\frac{1}{\Gamma_{0}+\Gamma_{1}}\right)}{r}\right]}{r+\Gamma_{0}+\Gamma_{1}},$ (28)
from which one can see that the critical time wouldn’t be affected by the coherence terms of the initial state.