### v3

# Accelerating Quantum Relaxation via Temporary Reset: A Mpemba-Inspired Approach

Abstract
Slow relaxation processes spanning widely separated timescales pose fundamental challenges for probing steady-state properties and engineering functional quantum systems, such as quantum heat engines and quantum computing devices. We introduce a protocol that enables exponential acceleration of relaxation in general Markovian open quantum systems by temporarily coupling the system to a reset channel, inspired by the quantum Mpemba effect. Crucially, this acceleration persists even when the slowest decaying Lindbladian modes form complex-conjugate pairs. Unlike previous approaches, which typically target a single mode, our protocol eliminates multiple relaxation modes simultaneously. Specifically, using a simple depolarizing channel, we show that the relaxation of a dissipative transverse-field Ising chain can be significantly accelerated, underscoring the experimental feasibility of the protocol. This framework provides a versatile tool for controlling quantum relaxation timescales, with broad implications for quantum thermodynamics, computation, and state preparation.
Introduction.—Relaxation processes in open quantum systems, wherein a driven or interacting system settles into an equilibrium or a nonequilibrium steady state, are of fundamental importance in nonequilibrium physics and for practical quantum technologies. The timescales of such relaxations often determine the feasibility and performance of quantum devices. For example, in quantum thermodynamics, the power output of heat engines or transport junctions is limited by how rapidly the working substance reaches its steady state in each cycle [1, 2, 3]. In addition, fast relaxation facilitates ground state laser cooling [4, 5, 6] and the reliable preparation of quantum states [7]. Faster relaxation can also lead to more efficient quantum algorithms [8], particularly in dissipative computational architectures where the output is encoded in the system’s stationary state [8, 9]. Slow relaxation inevitably allows other unwanted dissipative dynamics to consume substantial time resources, compromising the efficiency of the final stabilized state [10, 11]. In contrast, slow relaxation is sometimes desirable—such as when the computational task targets metastable states [12, 13, 14, 15], whose prolonged lifetimes are essential for robust information storage or approximate optimization.
Researchers have developed various techniques to accelerate relaxation toward both thermal equilibrium [16] and nonequilibrium steady states [17, 18, 11]. Most of these methods, however, are restricted to specific systems and rely on intricate external controls. For instance, some approaches require auxiliary Hamiltonians or full knowledge of the system Hamiltonian, limiting their applicability to simple setups. A more general and elegant strategy is to accelerate relaxation by tailoring the initial state distribution, without the need for continuous external driving. A particularly intriguing example of this approach is inspired by the so-called Mpemba effect [19], a counterintuitive relaxation anomaly in which a hotter system can cool faster than a colder one. This phenomenon implies the existence of an optimal initial state that minimizes the relaxation timescale. Theoretical frameworks for identifying such optimal states have been proposed in both classical [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and quantum systems [10, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], see also [43, 44] for recent reviews. However, existing quantum implementations typically impose strict constraints, such as assuming pure initial states or requiring the second-largest eigenvalue of the Lindbladian to be real. Moreover, careful initial-state design is a delicate task, since it relies on detailed knowledge of the original initial state and the system dynamics, which is typically not known a prior in complex quantum systems. These issues limit their applicability in general settings. Thus, a general and practical method for inducing faster relaxation in open quantum systems, which is robust with respect to different initial states, remains elusive.
In this study, we address these challenges by proposing an experimentally feasible protocol that can both accelerate and decelerate general quantum relaxation processes from any initial state, by resetting the system to a specified (possibly mixed) target state [45, 46, 47]. Here, this type of quantum reset operation can be realized through engineered dissipation, i.e., quantum reservoir engineering [48, 49, 9, 50], and should be distinguished from the conventional reset typically used for qubit initialization in quantum computation [51, 52, 53, 54, 55, 56, 57, 58]. In our protocol, we introduce a finite-duration reset phase into the open-system dynamics by temporarily coupling the system to a reset channel. We show that this protocol can exponentially accelerate relaxation, even when starting from mixed initial states and in the presence of complex decay modes. Remarkably, our method allows for the simultaneous elimination of multiple relaxation modes—a capability absent in previous approaches. Moreover, selective application of the protocol enables the system of interest to remain in metastable states for extended periods, offering a systematic route to suppress relaxation.
Setup.—We consider a general open quantum system defined on a Hilbert space $\mathcal{H}$ of dimension $d$, whose dynamics in the interaction picture are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) master equation $d\rho/dt=\mathcal{L}(\rho),$ where the Lindbladian superoperator $\mathcal{L}$ is given by
$\mathcal{L}(\rho)\equiv-i[H,\rho]+\sum_{i}\left[J_{i}\rho J_{i}^{\dagger}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},\rho\}\right].$ (1)
Here, $H$ is the Hamiltonian of the system and the jump operators $J_{i}$ describe the dissipative coupling to the environment. The evolution can be analyzed via the spectral decomposition of $\mathcal{L}$, whose eigenvalues are denoted $\lambda_{k}$ and ordered such that $0\geq\text{Re}(\lambda_{k})\geq\text{Re}(\lambda_{k+1})$. Assuming that $\mathcal{L}$ is diagonalizable, let $R_{k}$ and $L_{k}$ denote the corresponding right and left eigenmatrices, satisfying $\mathcal{L}(R_{k})=\lambda_{k}R_{k}$ and $\mathcal{L^{\dagger}}(L_{k})=\lambda_{k}^{\star}L_{k},\ k=1,...,d^{2}$, with biorthogonal normalization
$\text{Tr}(L_{k}^{\dagger}R_{h})=\delta_{kh}.$ (2)
The dual superoperator $\mathcal{L}^{\dagger}$ governs the Heisenberg-picture dynamics of observables:
$\mathcal{L}^{\dagger}(O)=i[H,O]+\sum_{i}\left[J_{i}^{\dagger}OJ_{i}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},O\}\right].$
Given an initial state $\rho_{0}$, the system state at time $t$ evolves as
$\rho(t)=e^{t\mathcal{L}}[\rho_{0}]=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}e^{\lambda_{k}t}c_{k}R_{k},$ (3)
where $c_{k}\equiv\text{Tr}(L_{k}^{\dagger}\rho_{0})$. The unique stationary state of the open quantum system $\rho_{\rm ss}$ is given by $\rho_{\rm ss}=\lim_{t\rightarrow\infty}\rho(t)=R_{1}$, assuming $\lambda_{1}=0$ is non-degenerate. In the long-time limit, the relaxation is dominated by the slowest decaying mode, and the deviation from stationarity obeys $||\rho(t)-\rho_{\rm ss}||\sim\exp\left(-|\text{Re}\lambda_{2}|t\right),$ so the characteristic relaxation timescale is given by $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$. For clarity of exposition, we first consider the simple case where the second eigenvalue $\lambda_{2}$ is real and non-degenerate, as in [10]. Later, we will generalize the discussion to include cases where the subleading eigenvalues form complex conjugate pairs.
Reset protocol to accelerate relaxation processes.—Recognizing that the slowest decaying mode governs the relaxation timescale, we propose a quantum reset protocol that suppresses its excitation from a general initial state via a finite-duration reset phase, thereby enabling exponential acceleration of relaxation. Specifically, we let the system evolve under modified dynamics for a finite time $t_{s}$, during which it is stochastically reset to a chosen state $\rho_{\delta}$ at random times, with events following a Poisson process of rate $r$. After time $t_{s}$, the reset channel is turned off, and the system resumes evolving under the original Lindbladian $\mathcal{L}$. That is, the system density matrix $\rho(t)$ evolves under the reset channel for $t\in[0,t_{s}]$ and decouples from the channel for $t>t_{s}$. The modified dynamics during the reset phase is governed by a new Lindbladian $\mathcal{L}_{r}$ [45], defined as
$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)\rho_{\delta}-r\rho.$ (4)
Expressing $\rho_{\delta}=\sum_{\alpha}p_{\alpha}|\psi_{\alpha}\rangle\langle\psi_{\alpha}|$, $\mathcal{L}_{r}$ is equivalent to introducing jump operators $J_{i}^{r,\alpha}=\sqrt{rp_{\alpha}}|\psi_{\alpha}\rangle\langle\phi_{i}|$ into $\mathcal{L}$, where $|\phi_{i}\rangle$ form an auxiliary complete orthonormal basis (see End Matter). The term $r\text{Tr}(\rho)\rho_{\delta}$ contributes only to the stationary mode and vanishes on all traceless eigenmodes $R_{i}$ with $i\geq 2$, which satisfy $\mathrm{Tr}(R_{i})=0$ due to $\mathrm{Tr}(L_{1}^{\dagger}R_{i})=0$ and $L_{1}^{\dagger}=\mathbb{I}$ [cf. (2)]. Consequently, the modified Lindbladian acts as
$\mathcal{L}_{r}(R_{i})=(\lambda_{i}-r)R_{i},\ i\in\{2,...,d^{2}\}.$ (5)
This induces a uniform spectral shift by $-r$ for all non-stationary eigenmodes, leaving the eigenmatrices ($R_{i}$, $i\geq 2$) themselves unchanged. In contrast, the stationary mode $R_{1}$ retains eigenvalue zero but is modified under reset. Let $\rho^{r}(t):=e^{t\mathcal{L}_{r}}\rho_{0}$ be the state at time $t$ under reset dynamics. Its spectral decomposition reads:
$\rho^{r}(t)$ $=e^{t\mathcal{L}_{r}}\rho_{\rm ss}+\sum_{k=2}^{d^{2}}c_{k}e^{(\lambda_{k}-r)t}R_{k}$
$:=\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}R_{k},$ (6)
where $c_{k}^{r}(t)$ are time-dependent mode amplitudes. Explicitly, these modified coefficients are given by (End Matter)
$c_{k}^{r}(t)\equiv\left[c_{k}-\frac{r\cdot d_{k}}{r-\lambda_{k}}\right]e^{-rt}+\frac{r\cdot d_{k}}{r-\lambda_{k}}e^{-\lambda_{k}t},$ (7)
where $d_{k}=\text{Tr}(L_{k}^{\dagger}\rho_{\delta})$. By properly choosing an optimal $t^{*}_{s}$, we can set $c_{2}^{r}(t^{*}_{s})=0$, thereby eliminating the overlap with the slowest decaying mode. Solving $c_{2}^{r}(t_{s})=0$ yields
$t^{*}_{s}(r)=\frac{1}{r-\lambda_{2}}\ln\left[1-\frac{c_{2}(r-\lambda_{2})}{r\cdot d_{2}}\right].$ (8)
By an appropriate choice of $\rho_{\delta}$ such that $c_{2}/d_{2}\leq 0$, $t^{*}_{s}(r)\geq 0$ can be assured. After the reset phase ($t>t_{s}$), the system returns to original Lindbladian evolution, now without the contribution from the slowest mode. The full evolution reads:
$\rho^{r}(t)=$ $\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t)R_{k}$
$+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}\Theta(t-t_{s})R_{k},$ (9)
where $\Theta(x)$ is the Heavside step function. This protocol with an optimal $t_{s}=t^{*}_{s}$ reduces the relaxation time to
$\tau_{{\rm rel}}\sim t^{*}_{s}+\frac{1}{|{\rm Re}(\lambda_{3})|},$ (10)
which is generically shorter than the unperturbed timescale $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$. Eq. (9) can be generalized to the dynamics of observables [59], which may be more convenient to measure.
Our protocol also enables controlled deceleration of relaxation. By choosing $\rho_{\delta}$ and $t_{s}$ to amplify the magnitude of $c_{k}^{r}(t)$, one can prolong the time during which the system remains away from the steady state. This capability could be useful in quantum metrology and computation, where long-lived metastable states play a functional role. Our reset protocol thus provides a general framework for controlling relaxation processes in open quantum systems.
Extension to complex-conjugate relaxation modes.—We now consider the more general case where the dominant relaxation modes form a complex-conjugate pair. In this case, the long-time dynamics of the density matrix takes the form
$\rho(t)=\rho_{\rm ss}+e^{\lambda_{2}t}c_{2}R_{2}+e^{\lambda_{2}^{\dagger}t}c_{2}^{*}R_{2}^{\dagger}+\mathcal{O}(e^{\lambda_{3}t}),$ (11)
where $c_{2}=\text{Tr}(L_{2}^{\dagger}\rho_{0})$ and $c_{2}^{*}=\text{Tr}(L_{2}\rho_{0})$ is its complex conjugate. Most known strategies are unable to simultaneously suppress both components of such complex relaxation modes.
Our reset protocol, however, naturally addresses this limitation. The modified coefficients under quantum reset are given by
$c_{2}^{r}(t)=\left[c_{2}-\frac{r\cdot d_{2}}{r-\lambda_{2}}\right]e^{-rt}+\frac{r\cdot d_{2}}{r-\lambda_{2}}e^{-\lambda_{2}t}$ , (12a)
$c_{2}^{r,*}(t)=\left[c_{2}^{*}-\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}\right]e^{-rt}+\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}e^{-\lambda_{2}^{*}t}$ (12b)
Importantly, we observe that these coefficients remain complex conjugates of each other for all $t$, i.e., $[c_{2}^{r}(t)]^{*}=c_{2}^{r,*}(t)$. Therefore, eliminating the entire complex mode reduces to a single complex condition:
$c_{2}^{r}(t_{s})=0\Leftrightarrow\text{Re}[c_{2}^{r}(t_{s})]=\text{Im}[c_{2}^{r}(t_{s})]=0,$ (13)
because $c_{2}^{r}(t_{s})=0\Rightarrow c_{2}^{r,*}(t_{s})=[c_{2}^{r}(t_{s})]^{*}=0$. In some cases, this condition can be exactly satisfied by tuning $r$ and $t_{s}$. Although Eq. (13) does not always admit an exact solution, in practice, it suffices to reduce $|c_{2}^{r}(t_{s})|$ to an exponentially small value ($\text{Re}[c_{2}^{r}(t_{s})]\ll 1,\ \text{Im}[c_{2}^{r}(t_{s})]\ll 1$), which still leads to exponential acceleration of relaxation. Additionally, we show in [59] that, if the condition
${\rm Re}(c_{2}^{*}d_{2})\leq|c_{2}|^{2}$ (14)
is satisfied, $|c_{2}^{r}(t_{s})|<|c_{2}|$ can always be achieved for any positive $t_{s}$ that below a certain threshold. This is a weak condition that be readily satisfied for many choices of $\rho_{\delta}$. It is strictly weaker than the commonly satisfied condition $c_{2}/d_{2}\leq 0$, which, in the case where $\lambda_{2}$ is real, guarantees the existence of a $t_{s}^{*}$ such that $c^{r}_{2}(t_{s}^{*})=0$. Eq. (14) thus provides a sufficient condition for reset-induced Mpemba-type acceleration [20]. It implies that fine-tuning of either $\rho_{\delta}$ or $t_{s}$ is unnecessary for achieving acceleration. We illustrate this behavior concretely in the following examples.
Example 1.—To illustrate our protocol, we first consider a minimal example: a two-level quantum system, which may represent a qubit or a spin-$1/2$ particle. While minimal, this example is crucial in many practical settings, as qubit reset often constitutes the bottleneck in large-scale quantum processes [60, 58]. We set the ground state energy $E_{0}=0$ and excited state energy $E_{1}=E$, such that the system Hamiltonian reads $H=E|1\rangle\langle 1|+\Omega(\sigma^{+}+\sigma^{-})$, where $\Omega$ is the intrinsic coherent coupling between the two levels, $\sigma^{+}=|1\rangle\langle 0|$ and $\sigma^{-}=|0\rangle\langle 1|$. The system is coupled to a thermal environment through the Lindblad jump operators $J_{0}=\sqrt{\gamma_{0}}\sigma^{+}$ and $J_{1}=\sqrt{\gamma_{1}}\sigma_{-}$. Here, $\gamma_{0}$ ($\gamma_{1}$) is the transition rate from state $|0\rangle$ to state $|1\rangle$ (or converse). Assuming the rates satisfy detailed balance with inverse temperature $\gamma_{0}e^{-\beta_{\text{env}}E_{0}}=\gamma_{1}e^{-\beta_{\text{env}}E_{1}}$, i.e., $\gamma_{0}=\gamma_{1}e^{-\beta_{\text{env}}E},$ the system evolves to a unique equilibrium state $\rho_{\rm eq}=\frac{e^{-\beta_{\text{env}}H}}{\text{Tr}(e^{-\beta_{\text{env}}H})}$ irrespective of the initial condition. We initialize the system in a general mixed state:
$\hat{\rho}_{0}$ $=\left(\begin{array}[]{cc}\frac{1}{1+e^{-\beta_{0}E}}&ke^{i\phi}\\ ke^{-i\phi}&\frac{e^{-\beta_{0}E}}{1+e^{-\beta_{0}E}}\end{array}\right),$
$0\leq k$ $\leq\frac{\sqrt{e^{\beta_{0}E}}}{e^{\beta_{0}E}+1},\quad 0\leq\phi\leq 2\pi,$
where the off-diagonal terms $ke^{\pm i\phi}$ quantify initial quantum coherence. In the absence of intrinsic transitions ($\Omega=0$), the Hamiltonian is diagonal and the relaxation process reduces to pure decoherence. In this case, all eigenmodes and eigenvalues of the Lindbladian can be computed analytically [59]. When we apply our reset protocol with the reset state $\rho_{\delta}=|0\rangle\langle 0|$, corresponding to the ground state to ensure $t_{s}(r)\geq 0$, the two slowest decaying modes form a complex-conjugate pair:
$c_{2}^{r}(t)=e^{-rt+i\phi}k,\quad c_{2}^{r,*}(t)=e^{-rt-i\phi}k.$ (15)
This overlap cannot be entirely eliminated in finite time, but it decays exponentially fast, leading to a relaxation timescale
$\tau_{{\rm rel}}\sim\frac{1}{r}+\frac{1}{|\text{Re}(\lambda_{3})|},$ (16)
where $1/r$ characterizes the reset protocol timescale, and $1/|\text{Re}(\lambda_{3})|$ represents the accelerated relaxation timescale with the slowest mode eliminated. Clearly, $\tau_{{\rm rel}}<1/|{\rm Re}(\lambda_{2})|$ for sufficiently large $r$. By appropriately tuning the reset rate, the relaxation timescale can approach $1/|\rm Re(\lambda_{3})|$ arbitrarily closely. Using Eq. (9), the full dynamics under our reset protocol can thus be determined.
Since the condition $c_{2}^{r}(t_{s})=0$ cannot be satisfied exactly in this case—though $|c_{2}^{r}(t_{s})|$ can be made exponentially small by choosing $t_{s}\sim 1/r$—we instead define $t_{s}$ such that a faster decaying mode is suppressed. Specifically, solving $c_{4}^{r}(t_{s})=0$ gives
$t_{s}=\frac{\ln\left[1+\frac{\gamma_{1}k_{\text{tot}}}{r}\left(\frac{1}{\gamma_{0}+\gamma_{1}e^{(\beta_{0}-\beta_{\text{env}})E}}-\frac{1}{\gamma_{0}+\gamma_{1}}\right)\right]}{k_{\text{tot}}},$ (17)
where the total transition rate is $k_{\text{tot}}=r+\gamma_{0}+\gamma_{1}$. This choice fully eliminates the fastest decaying relaxation mode. In addition, for a sufficiently large reset rate $r$, the slowest mode’s amplitude becomes exponentially small after the reset phase. Alternative selections for $t_{s}$ can also achieve accelerated relaxation, indicating that detailed dynamics knowledge is not essential. Notably, if the initial inverse temperature satisfies $\beta_{0}<\beta_{\text{env}}$, a scenario analogous to “cooling”, the ground state $|0\rangle\langle 0|$ serves as a useful $\rho_{\delta}$ regardless of the rate $r$. Conversely, when $\beta_{0}>\beta_{\text{env}}$, resetting to the ground state decelerates relaxation, necessitating the choice $\rho_{\delta}=|1\rangle\langle 1|$ for acceleration. This example demonstrates that, by appropriately choosing $\rho_{\delta}$, one can switch between reset-induced acceleration and deceleration of relaxation.
We next present numerical results for illustration. We characterize the distance between the transient state $\rho(t)$ and the stationary state $\rho_{\rm ss}$ using the standard trace distance $D[\rho(t)|\rho_{\rm ss}]:=\text{Tr}|\rho(t)-\rho_{\rm ss}|/2$ where $|A|:=\sqrt{A^{\dagger}A}$. Other measures, such as the $L_{\infty}$ norm $D_{\infty}[\rho(t)|\rho_{\rm ss}]:=\max_{i}|\lambda_{i}|$ [$\lambda_{i}$ is the $i$-th eigenvalue of $\rho(t)-\rho_{\rm ss}$], yield qualitatively similar outcomes. We first plot $D[\rho(t)|\rho_{\rm eq}]$ over $t$ for various rate $r$ in Fig. 1 (a). The parameters for $r>0$ cases are: initial conditions $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$; The $r=0$ scenario has $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$. The environment is at lower temperature $\beta_{\text{env}}=4.0$. Although the $r>0$ cases are initially farther from equilibrium compared to the $r=0$ case, the reset protocol substantially accelerates relaxation so that the former cases reach equilibrium even faster. We next plot $t_{s}$ as a function of $r$ in Fig. 1 (b) and (d), which show that $t_{s}$ decreases monotonically as $r$ increases. Fig. 1 (c) confirms accelerated relaxation for $\Omega=2$, here measured by the $L_{\infty}$ norm to reduce fluctuations in the curves.
Figure 1: Relaxation dynamics of the two-state system. (a) and (c): Distances between $\rho(t)$ and $\rho_{\text{eq}}$ as the function of time $t$ with different resetting rate $r$. (b) and (d): The $t_{s}$ as a function of the resetting rate $r$. The parameters are chosen as $\beta_{0}=2.0$, $k=0.32$ and $\phi=1$ with reset protocol and $\beta_{0}=3.0,$ $k=0.21$ and $\phi=1$ without reset. For (a) and (b) $\Omega=0$, for (c) and (d) $\Omega=2.0$.
Example 2.—To illustrate the potential feasibility of our protocol in complex quantum systems, we consider another example, the dissipative transverse-field Ising model (TFIM), a paradigmatic system relevant to various quantum platforms. We numerically study an open dissipative TFIM of length $N$ with Hamiltonian
$H=-J\sum_{i=1}^{N-1}\sigma_{i}^{z}\sigma_{i+1}^{z}-g\sum_{i=1}^{N}\sigma_{i}^{x},$ (18)
and jump operators $J_{i,\downarrow}=\sqrt{\gamma}\,\sigma_{i}^{-},\quad J_{i,\uparrow}=\sqrt{\gamma e^{-\beta}}\;\sigma_{i}^{+}$. Here, $\sigma^{x,z}_{i}$ are Pauli operators, $\sigma_{i}^{-}=|0\rangle\langle 1|$ and $\sigma_{i}^{+}=|1\rangle\langle 0|$. We fix $J=1$, $g=1.2$, $\gamma=0.5$, $\beta=1/k_{B}T=1$ throughout. We choose $\rho_{\delta}=\mathbb{I}/d$ (with $d=2^{N}$). The initial state $\rho_{0}$ is chosen by normalizing $\rho_{\rm ss}+\alpha V_{2}$, with $\rho_{\rm ss}$ being the steady-state and $V_{2}$ being the second normalized eigenvector of the Lindbladian. A larger $\alpha$ implies that $\rho_{0}$ is farther from $\rho_{\rm ss}$. Here, $\rho_{\rm ss}$ is nontrivial, with nonzero entanglement. Parameters: $t_{s}=0.5\,\tau_{2}$ and $t_{s}=0.08\,\tau_{2}$ with $\tau_{2}=1/|\operatorname{Re}\lambda_{2}|$, the relaxation timescale, total duration $T=6\,\tau_{2}$, $\alpha=0.55/0.05$. Notably, $t_{s}$ is chosen arbitrarily, rather than being determined a prior. The accelerated relaxation processes under different reset rate are shown in Fig. 2. We also examine cases with different parameters in [59], where acceleration is consistently observed. In Fig. 2 (b) and (d), genuine Mpemba effects are observed, where the system is first driven away from the steady state by the reset protocol compared to the no-reset case, and then reaches stationarity faster. The existence of such an effect broadens the range of usable reset states.
Some remarks are in order regarding experimental implications. In general, the optimal duration $t_{s}$ of the reset phase cannot be determined a priori from theoretical considerations, as the relevant dynamical details are typically unknown in complex quantum systems. The TFIM example shows how this challenge can be overcome: fine-tuning of $t_{s}$ is not required—any finite value shorter than the native relaxation time leads to substantial speed-up, and $t_{s}$ can be further adjusted a posteriori based on observed dynamics. This robustness, without requiring prior knowledge of the system dynamics and initial states, distinguishes our protocol from previous approaches. Additionally, the reset channel in the example corresponds to a depolarizing channel, which is conceptually straightforward to implement, e.g., by applying an isotropic white-noise field or coupling an infinite-temperature reservoir.
Figure 2: Acceleration of the relaxation of TFIMs. Reset rate $r$ are chosen as $0,1.0,5.0,10.0,20.0$. (a) $N=5$ ($d=32$), $\alpha=0.55$, $t_{s}=0.08\tau_{2}$. (b) $N=5$, $\alpha=0.05$, $t_{s}=0.50\tau_{2}$. (c) $N=6$ ($d=64$), $\alpha=0.55$, $t_{s}=0.08\tau_{2}$. (d) $N=6$, $\alpha=0.05$, $t_{s}=0.50\tau_{2}$. The vertical gray lines indicate $t_{s}$ when the reset channel is removed. Inset: Schematic representation of the dissipative TFIM, with each spin coupled to a thermal bath.
Experimental feasibility in general cases.—The TFIM example illustrates the power of our protocol: the acceleration of relaxation via a depolarizing channel has the potential to be applied to complex quantum platforms. However, in general, it may be necessary to choose other $\rho_{\delta}$ to achieve acceleration. For few-body systems (such as single Rydberg atoms or other multi-level emitters), the protocol may be implemented with quantum reservoir engineering. Such techniques have become increasingly routine for tailoring dissipative dynamics in diverse quantum systems, including atoms [61, 62, 63], trapped ions [48, 64], superconducting circuits [61, 50, 65], and optomechanical setups [66, 67, 68]. For quantum many-body systems, an exact implementation via reservoir engineering generally requires high-order interactions. Such engineered $N$-body dissipators have been put forward theoretically [69, 70] and realized experimentally up to four-body interactions [71].
A complementary route is to approximate the reset channel: combine any available state-preparation protocol for $\rho_{\delta}$, whether reservoir engineering, measurement-based feedback, or other methods, with Trotterization [72, 73, 74] (see End Matter). This strategy sidesteps the need to engineer reset jump operators directly and therefore avoids high-order couplings even in many-body systems. The main experimental challenge then reduces to preparing $\rho_{\delta}$. Encouragingly, fast and high-fidelity state preparation has been reported across diverse platforms [75, 76, 63, 77]. Notably, Ref. [77] realized state preparation in a 35-spin TFIM with Trotterization. This suggests that our protocol may be testable and applicable on existing platforms.
Concluding remarks.—We have introduced a general framework for accelerating relaxation in open quantum systems via temporary reset, applicable to arbitrary initial states. The proposed protocol can suppress multiple relaxation modes simultaneously, enabling enhanced control over the relaxation dynamics. As a practical example, we demonstrated that the relaxation of a TFIM can be significantly accelerated by using a simple depolarizing channel. This example highlights a powerful feature of our protocol: leveraging easily prepared states to accelerate the preparation of states that are difficult to reach. Furthermore, introducing temporary dephasing noise—a different type of channel—can also accelerate relaxation (End Matter). This suggests the broader applicability of our central idea: temporarily coupling the system to various quantum channels may provide a general route to enhanced relaxation. In future studies, our approach could be extended to certain non-Markovian and Floquet dissipative systems (e.g., using ideas in [10, 78]). Overall, our results establish temporary reset as a powerful and experimentally feasible tool for controlling relaxation timescales in open quantum dynamics.
This work is partly supported by the Innovation Program for Quantum Science and Technology (2021ZD0303306), MOST(2022YFA1303100).
End Matter
Derivation of the modified Lindbladian $\mathcal{L}_{r}$.—Here, we derive Eq. (4) of the main text. Let $\{|\phi_{i}\rangle\}_{i=1}^{d}$ be an orthonormal basis, satisfying $\sum_{i}|\phi_{i}\rangle\!\langle\phi_{i}|=\mathbb{I}$. The reset state is generally written as $\rho_{\delta}=\sum_{\alpha}p_{\alpha}|\psi_{\alpha}\rangle\!\langle\psi_{\alpha}|$. Introduce jump operators
$J_{i,\alpha}^{r}=\sqrt{r\,p_{\alpha}}|\psi_{\alpha}\rangle\!\langle\phi_{i}|,\quad i=1,\dots,d.$
The associated Lindblad term is
$\mathcal{D}_{r}(\rho)=\sum_{i=1}^{d}\!\bigl{(}J_{i}^{r}\rho\,J_{i}^{r\dagger}-\tfrac{1}{2}\{J_{i}^{r\dagger}J_{i}^{r},\rho\}\bigr{)}.$ (19)
Direct evaluation gives
$\sum_{i,\alpha}J_{i,\alpha}^{r}\rho\,J_{i,\alpha}^{r\dagger}=r\,\rho_{\delta}\,\text{Tr}[\rho]$
$-\tfrac{1}{2}\sum_{i,\alpha}\{J_{i,\alpha}^{r\dagger}J_{i,\alpha}^{r},\rho\}=-r\,\rho,$
where $\text{Tr}[\rho]=\sum_{i}\langle\phi_{i}|\rho|\phi_{i}\rangle$, $\sum_{i}|\phi_{i}\rangle\!\langle\phi_{i}|=\mathbb{I}$ and $\langle\psi_{\alpha}|\psi_{\alpha}\rangle=1$ have been used. This yields $\mathcal{D}_{r}(\rho)=r\bigl{[}\text{Tr}[\rho]\,\rho_{\delta}-\rho\bigr{]}$. Hence, we arrive at Eq. (4),
$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}[\rho]\,\rho_{\delta}-r\rho.$ (20)
Derivation of the modified coefficients.—Here, we derive Eq. (7), the explicit form of the modified coefficients with reset. Acting the modified semi-group $e^{t\mathcal{L}_{r}}$ on both sides of
$\rho_{0}=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})R_{k}$ (21)
yields
$\rho^{r}(t)=e^{t\mathcal{L}_{r}}\rho_{0}=e^{t\mathcal{L}_{r}}\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})e^{(\lambda_{k}-r)t}R_{k}.$ (22)
To proceed, we use an explicit form of the steady state under reset dynamics [45]
$\rho_{\rm ss}^{r}=\lim_{t\rightarrow\infty}\rho^{r}(t)=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\frac{rd_{k}}{r-\lambda_{k}}R_{k}.$ (23)
Then, expressing the steady-state condition $e^{t\mathcal{L}_{r}}\rho_{\rm ss}^{r}=\rho_{\rm ss}^{r}$ with Eq. (23) and applying Eq. (5), we get
$e^{t\mathcal{L}_{r}}\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\frac{rd_{k}}{r-\lambda_{k}}e^{(\lambda_{k}-r)t}R_{k}=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\frac{rd_{k}}{r-\lambda_{k}}R_{k},$
from which we solve $e^{t\mathcal{L}_{r}}\rho_{\rm ss}$. Substituting $e^{t\mathcal{L}_{r}}\rho_{\rm ss}$ into Eq. (22) and comparing with Eq. (Accelerating Quantum Relaxation via Temporary Reset: A Mpemba-Inspired Approach), we identify modified coefficients $c_{k}^{r}(t)$ defined in Eq. (7).
Realizing the reset protocol with arbitrary state preparation processes.—We now show that any state preparation protocol could approximately realize the reset channel in our scheme via a Trotterization-based construction. Notably, Trotterization has already been implemented experimentally in superconducting quantum circuits [79], IBM Quantum’s hardware [74] and even quantum many-body platform [77] to generate dissipative dynamics, such as the dissipative TFIM. Moreover, tensor network methods offer promising avenues for extending Trotterization to more complex open quantum systems [72]. We first consider the case where the original dynamics is absent [$\mathcal{L}(\rho)=0$] to gain insight. In this case, the generator corresponding to the reset channel simply reads
$\dot{\rho}(t)=\mathcal{R}[\rho(t)]=r[\rho_{\delta}-\rho(t)].$ (24)
Thus, a direct integration shows that applying the reset channel for a time $t_{s}$ maps an initial state $\rho_{0}$ to a final state $\rho(t_{s})$ according to
$\rho(t_{s})=e^{\mathcal{R}t_{s}}[\rho_{0}]=p\rho_{0}+(1-p)\rho_{\delta},\quad p:=e^{-rt_{s}}.$ (25)
That is, a reset channel with rate $r=-\ln p/t_{s}$ is effectively equivalent to a state preparation protocol for $\rho_{\delta}$, with success probability $1-p$. $p$ is a classical probability generated beforehand. One simply flips a biased coin (or, equivalently, performs any local measurement that yields the desired classical randomness): with probability $1-p$ the system is driven into $\rho_{\delta}$ by any suitable operations and with the probability $p$ it is left untouched. Since there are no restrictions on the preparation method, once $\rho_{\delta}$ is locally preparable (e.g. a separable state), this operation can be implemented purely with local operations, even for many-body systems. Specifically, when $\rho_{\delta}=\mathbb{I}/d$, Eq. (25) is a depolarizing channel, which can be directly realized by the 2-design [80] without flipping a coin.
When $\mathcal{L}\neq 0$, we apply the Lie-Trotter formula [81]
$e^{(\mathcal{L}+\mathcal{R})t}=\lim_{n\to\infty}(e^{\mathcal{R}\frac{t}{n}}e^{\mathcal{L}\frac{t}{n}})^{n}.$ (26)
to realize the protocol. With this formula, the reset protocol during $[0,t_{s}]$ can be approximated as
$e^{(\mathcal{L}+\mathcal{R})t_{s}}[\rho_{0}]\approx(e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t})^{n}[\rho_{0}],$ (27)
where $\delta t:=t_{s}/n\ll 1$. Using Eq. (25), we have
$e^{\mathcal{R}\delta t}[\rho]\approx(1-p_{s})\rho+p_{s}\rho_{\delta},$ (28)
which can be interpreted as a state preparation mapping from $\rho$ to $\rho_{\delta}$, with success probability $p_{s}:=r\delta t$. Practically, $r$ is determined by $p_{s}$ and $\delta t$, whose values are set by experimentalists.
Thus, the reset protocol may be achieved experimentally by performing the aforementioned state preparation operations stroboscopically: at each discrete time point $t_{i}=i\delta t,\ (i=1,...,n)$ within $[0,t_{s}]$, one applies the state preparation mapping from $\rho(t_{i})$ to $\rho_{\delta}$ with a given probability $p_{s}$. There are no constraints on the preparation time $\tau_{\rm prep}$ for $\rho_{\delta}$, but on average it adds a cost of $np_{s}\tau_{\rm prep}$ to the tailored relaxation timescale. Therefore, it is desirable to choose a $\rho_{\delta}$ that can be prepared efficiently, e.g. the maximally mixed state in our TFIM example. This state is convenient to prepare: coupling to an infinite-temperature bath can achieve it in $\mathcal{O}(\ln N)$ time [82] with minimal resources in a $N$-body system—typically negligible compared to the system’s intrinsic relaxation time. Local control methods could in principle realize it in constant time, though with resources scaling with $N$.
This Trotterization-based method admits a natural interpretation at the level of stochastic trajectories: for a Poissonian reset process with rate $r$, the system is stochastically reset to the target state with probability $r\delta t$, in each small interval $\delta t$.
The approximation error is of order $\mathcal{O}(t_{s}^{2}/n)$ [83, 84, 73]. Explicitly, we establish a rigorous upper bound [59]:
$||e^{(\mathcal{L}+\mathcal{R})t_{s}}-(e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t})^{n}||\leq\frac{t_{s}^{2}}{2n}||[\mathcal{L,R}]||,$ (29)
where |||||$\cdot$||| | ⋅ | | denotes any norm that is contractive under Lindbladian evolution, such as the trace norm or the diamond norm. This bound holds for arbitrary Lindbladian $\mathcal{L}$ and $\mathcal{R}$. In our specific case,
$[\mathcal{L,R}]\rho:=[\mathcal{LR}-\mathcal{RL}]\rho=r[\mathcal{L}(\rho_{\delta})-\rho_{\delta}].$ (30)
Hence, a large $n$ is not required when $t_{s}\ll 1$, $rt_{s}\ll 1$ or $\rho_{\delta}$ is close to $\rho_{\rm ss}$. A small $t_{s}$ with moderate $r$ offers a practical regime that may yield substantial acceleration and simultaneously reduce the required number of Trotter steps.
Additionally, higher-order approximations could be used to reduce the error for fixed $n$. For instance, the second-order Suzuki-Trotter formula $e^{(\mathcal{L}+\mathcal{R})t_{s}}\approx(e^{\mathcal{L}\delta t/2}e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t/2})^{n}$ suppresses the error to $\mathcal{O}(t_{s}^{3}/n^{2})$ [83, 73].
Dephasing noise can accelerate the relaxation of the TFIM.—Here, we add dephasing noise along a single axis (typically the $z$-axis) to the TFIM. The corresponding jump operators are $L^{(\phi)}_{i}=\sqrt{\gamma_{\phi}}\sigma_{i}^{z},\quad i=1,\dots,N$. This contributes to the total Lindbladian via an extra term
$\mathcal{L}_{\phi}[\rho]=\gamma_{\phi}\sum_{i=1}^{N}\left(\sigma_{i}^{z}\rho\sigma_{i}^{z}-\rho\right),$ (31)
which could be interpreted as a partial and local reset channel (only the $z$-axis is affected, and only local one-body jump operators are involoved). As shown in Fig. 3, the relaxation is accelerated significantly.
Figure 3: Dephasing noise induce acceleration of relaxation in a 5-site TFIM. Parameters: (a) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.2\tau_{2}$ (b) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.8\tau_{2}$.
A theoretical argument supporting the finding is provided in [59].